Pam and Jena, working together, can weed the garden in 9 hours. Working alone, Jena takes three times as long as Pam. How long does it take Pam to weed the garden alone?
Answer by Boreal(15235)

in 1 hour, both do 1/9 garden.
in 1 hour, Jena does 1/x garden.
in 1 hour, Pam does 3/x garden, 3 times as much.
(1/x)+(3/x)=(1/9)
multiply everything by 9x
9+27=x
x=36 hours.
Pam does (3/36) of garden (1/12) in an hour. It takes her 12 hours. ANSWER
Jena does (1/36) of the garden in an hour. It takes her 36 hours, 3 times as long
(1/12)+(1/36)=(4/36)=(1/9)